# Simulated Interactive Debugging

### v1

@@ 4.3. Results & Discussion
In total, all eight participants were able to solve the two programming tasks in the provided time. As shown in Figure 4, the most effective feature is the automatic breakpoint setting (5/8 agreement), followed by interactive debugging and chatting (4/8) and the explanation of automatic breakpoints (4/8). The features for progressive hints and test error explanation are only positively evaluated by 2/8 students. The progressive hints helped the students to identify where their code has a fault but did not reveal the solution. Instead, it attempted to guide the students to find the solution on their own to foster a learning effect. However, the participants perceived this feature as often too vague or too general. For most participants, the test error explanation helped to analyze the coding issue, while the feature itself could be made more interactively, e.g., allowing for different/alternative explanations. The explanation of the automatically set breakpoints helped most students to understand the wrong parts. Therefore, we conclude that it is a good addition to the automatic setting of breakpoints. The interactive debugging and chatting was able to add more detailed information to the debugging efforts. One participant mentioned that this even accelerated the overall debugging process. Others still found that the responses were vague and that more prompts could improve the feature. The automatic setting breakpoints feature was considered very effective as it set the breakpoints at the right places. One participant summarized its effectiveness as follows:
”It did get the breakpoint in the right place, which was a good start in debugging.”
One of the first hurdles of conventional interactive debugging is to set breakpoints. Our concept of simulated interactive debugging helps them to get started, and therefore, guides them along a deeper understanding of their coding errors and strives to develop debugging skills. The following quote from one of our participants about the interactive debugging and chatting highlights another aspect:
”It helped me brainstorm while guiding me and
not letting me fully rely on AI”
Our goal is to guide students to learn debugging but also the usage of AI-based tools like ChatGPT. By constructing a chatbot that uses LLMs, we can control the prompts and the contents of responses, and hence, still foster a learning experience for the students, e.g., by filtering direct solutions.
@@
Using the standardized system usability survey by Brooke (Brooke, 1996), we can conclude that overall, the participants are satisfied with the prototype’s usability. The average System Usability Scale (SUS) is around 65 (out of 100), providing a good basis for our future work. For example, 7/8 participants would like to use the system frequently, and 5/8 agreed that the system is easy to use. These findings confirm that our efforts are heading in the right direction but also leave space for further improvement.
@@
After completing the tasks, we also asked the participants how confident they are about setting breakpoints on their own. The responses indicated mixed confidence with a tendency to lower confidence: 1/8 participants strongly agreed to be confident, 2/8 agreed, 2/8 noted a neutral opinion, and 3/8 disagreed. Afterwards, we asked them to set breakpoints for a given code snippet (see Listing 1). The interesting lines are 6 to 7 because the handling of None values is missing.
Listing 1: Code snippet for setting a manual breakpoints.
1# Calculate the average of a list of
2# grades, if None, you should continue
3# to the next grade
4def calculate_average (grades):
5 total = 0
6 for grade in grades:
7 total += grade
8 average = total / len (grades)
9 return average
10
11# Test the function
12grades = [85, 90, 78, None, 92, None]
13average_grade = calculate_average(grades)
14print(f”Theaverageis:{average_grade}”)
6/8 participants would set breakpoints either at lines 6 or 7, or both, which shows that actually most of the participants have a correct intuition about where to set the breakpoint for further inspection. While we see good results for this relatively simple example, we conclude that due to overall low confidence, the automated setting of breakpoints would still be an interesting feature. However, such a feature could be applied only as hints or as confirmation for the students’ manually set breakpoints. Additionally, we asked the participants about their perceived key difference to their usual debugging process. The majority (6/8) mentioned that our simulated interactive debugging tool helped them to debug the coding issue and guided them systematically. However, one participant mentioned that interactive debugging could be more time-consuming. The following two quotes are representative responses:
”Normally I would just trial and error until I got it right, but this tool allowed me to systematically identify the problems in the code step by step.” ”My usual debugging process includes using print statements and doing everything myself but sometimes I just get tired and ask ChatGPT. Simulated Interactive Debugging helped me think while guiding me through the right process.”
@@
At the end of the post-task survey, we asked the participants about any additional features they would like to see in the Simulated Interactive Debugger tool that have not been implemented yet. 6/8 participants indicated that further hints about what and how to change would be helpful. For example, when the students really struggle, the chatbot could provide more detailed explanations and propose changes that the student could select. Additionally, one participant mentioned that the user interface could be improved to enable easier switching between the debugging information from the assisted debugger and the actual interactive debugging tool in the IDE. Lastly, one participant highlighted that integrating some automated test case generation would be helpful. For example, such test cases could be provided if the students cannot find an appropriate input for debugging.
4.4. Threats to Validity
Since we only performed a first pilot study with eight CS students, our results may not generalize. We focused our user study on a small number of students to receive focused feedback for the next development phase. Nevertheless, the results of our profiling survey showed that the participants are in our target, i.e., beginners in programming with no or only very limited experience with (interactive) debugging. For the next cycle of user studies, we will attempt to reach a larger set of participants. In order to mitigate the threat of manual errors in our analysis, we ensured that all results were analyzed and agreed upon by two authors independently.
5. Conclusion and Future Work
We proposed the concept of simulated interactive debugging as a first step to automate the teaching of debugging skills. Incorporating such techniques in the CS curriculum will be essential to teaching debugging beyond ad-hoc approaches and guiding students in using AI-based programming tools. In future, we plan to improve our interactive chatbot with more guidance, e.g., by using a state machine-driven approach, similar to Bouzenia et al. (Bouzenia et al., 2024). We also will explore the integration of several state-of-the-art APR techniques, implement the test assessment features for lecturers and students, and investigate how we can automatically help students to generate and improve their test suites.

### v2

@@ IV-B Results & Discussion
In total, all eight participants were able to solve the two programming tasks in the provided time. As shown in Figure 2, the most effective feature is the automatic breakpoint setting (5/8 agreement), followed by interactive debugging and chatting (4/8) and the explanation of automatic breakpoints (4/8). The features for progressive hints and test error explanation are only positively evaluated by 2/8 students. The progressive hints helped the students to identify where their code has a fault but did not reveal the solution. Instead, it tries to guide the students to find the solution on their own to foster a learning effect. However, the participants perceived this feature as often too vague or too general. For most participants, the test error explanation helped to analyze the coding issue, while the feature itself could be made more interactively, e.g., allowing for different/alternative explanations. The explanation of the automatically set breakpoints helped most students to understand their faults. Therefore, we conclude that it is a good addition to the automatic setting of breakpoints. The interactive debugging and chatting was able to add more detailed information to the debugging efforts. One participant mentioned that this even accelerated the overall debugging process. Others still found that the responses were vague and that more prompts could improve the feature. The automatic setting breakpoints feature was considered very effective as it set the breakpoints at the right places. One participant summarized its effectiveness as follows:
“It did get the breakpoint in the right place, which was a good start in debugging.”
One of the first hurdles of conventional interactive debugging is to set breakpoints. Our concept of simulated interactive debugging helps the students to get started, and therefore, guides them along a deeper understanding of their coding errors and strives to develop debugging skills. The following quote from one of our participants about the interactive debugging and chatting highlights another aspect:
“It helped me brainstorm while guiding me and
not letting me fully rely on A”
Our goal is to guide students to learn debugging but also the usage of AI-based tools like ChatGPT. By constructing an LLM-based chatbot, we can control the prompts and the content of responses, and hence, still foster a learning experience for the students, e.g., by filtering direct solutions.
@@
Using the standardized system usability survey by Brooke [33], we can conclude that overall, the participants are satisfied with the prototype’s usability. The average System Usability Scale (SUS) is around 65 (out of 100), providing a good basis for our future work. For example, 7/8 participants would like to use the system frequently, and 5/8 agreed that the system is easy to use.
@@
After completing the tasks, we also asked the participants how confident they are about setting breakpoints on their own. The responses indicated mixed confidence with a tendency to lower confidence: 1/8 participants strongly agreed to be confident, 2/8 agreed, 2/8 noted a neutral opinion, and 3/8 disagreed. Afterwards, we asked them to set breakpoints for a given buggy code snippet that calculates the average for a given list of grades without catering for None values in the list. 6/8 participants would set breakpoints at the correct lines, which shows that actually most of the participants have a correct intuition about where to set the breakpoint for further inspection. While we see good results for this relatively simple example, we conclude that due to overall low confidence, the automated setting of breakpoints would still be an interesting feature. However, such a feature could be applied only as hints or as confirmation for the students’ manually set breakpoints. Additionally, we asked the participants about their perceived key difference to their usual debugging process. The majority (6/8) mentioned that our Simulated Interactive Debugging tool helped them to debug the coding issue and guided them systematically. However, one of the participants mentioned that the interactive debugging process could be more time-consuming, though. The following two quotes are representative responses:
“Normally I would just trial and error until I got it right, but this tool allowed me to systematically identify the problems in the code step by step.” “My usual debugging process includes using print statements and doing everything myself but sometimes I just get tired and ask ChatGPT. Simulated Interactive Debugging helped me think while guiding me through the right process.”
@@
At the end of the post-task survey, we asked the participants about any additional features they would like to see in the Simulated Interactive Debugger. 6/8 participants indicated that further hints about what and how to change would be helpful. For example, when a student really struggles, the chatbot could provide more detailed explanations and propose changes that the student could select. One participant mentioned that the UI could be improved to allow easier switching between the assisted debugger and the interactive debugging tool in the IDE. Lastly, one participant highlighted that automated test generation would be helpful, e.g., if students cannot find an appropriate input for debugging.
IV-C Threats to Validity
We performed a first, small pilot study with eight CS students, i.e., our results may not generalize. We focused on a small number of students to receive focused feedback for the next development phase. The results of our profiling survey showed that the participants are in our target, i.e., beginners in programming with no or very limited (interactive) debugging experience. For the next cycle of user studies, we will try to reach a larger set of participants. To mitigate the threat of manual errors in our analysis, we ensured that all results were analyzed and agreed upon by two authors independently.
V Conclusion and Future Work
We proposed the concept of Simulated Interactive Debugging as a first step to automate the teaching of debugging skills. Incorporating such techniques in the CS curriculum will be essential to teach debugging beyond ad-hoc approaches and guide students in using AI-based programming tools. In future, we plan to incorporate more active guidance, e.g., by using a state machine-driven approach, similar to Bouzenia et al. [34]. We will also integrate state-of-the-art APR techniques and implement the test assessment and generation features.
Data Availability
Our artifact with the prototype and the study artifacts is available: https://doi.org/10.6084/m9.figshare.28202336
Acknowledgment
Part of this work was supported by the Ministry of Education, Singapore, under the Tertiary Research Fund (MOE-TRF) Grant No. MOE2023-TRF-034. We gratefully acknowledge this support, which made this research possible.